A piece of equipement will function only all components A, B and C are working. The probability of failing duirng one year is 15%, that of B failing in 5% & that of C failing in 10%. What is the probaibility that the equipment will fail before the end of the year?

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Shubham,

Ans:- The prob that the instrument will fail=1- prob that it would not fail

Now the prob that it would not fail=prob that the three parts ABC work together=(0.85)*(0.95)*(0.9)=0.72675

Hence the required prob is=1 - 0.72675=0.27325(ans) 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Shubham !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR